OpenStudy (anonymous):

stuck help!:) find the sum of 2sqrt6 and sqrt24

OpenStudy (accessdenied):

Hint: try finding perfect square factors in 24 and removing them by the rule \( \sqrt{ab} = \sqrt{a}\sqrt{b} \)

OpenStudy (anonymous):

what he said, multiply the square roots then square root the product and multiply by the coefficient

OpenStudy (lgbasallote):

my way to do this is to factor 24 into prime numbers so... \(\large \sqrt{24} = \sqrt{(2)(2)(2)(3)}\) take out of the sqrt all numbers that appeared twice

OpenStudy (anonymous):

yep did it and got 4 sqrt 6

OpenStudy (anonymous):

multiply whats is under the square root

OpenStudy (lgbasallote):

\(\Huge \checkmark\) @katlin95

OpenStudy (anonymous):

!:)